OpenStudy (angelina__nievse):

Can you help me convert 8.132132132 to a rational expression in the form of a/b were b = 0

OpenStudy (mjdennis):

Sure. Let's drop the "8" and find the fraction that matches 0.132... first, then you can put the "8" back in

OpenStudy (angelina__nievse):

132/99

OpenStudy (angelina__nievse):

Would that be the answer 8 132/99

OpenStudy (mjdennis):

Sorry, distracted.

OpenStudy (angelina__nievse):

Is that the answer

OpenStudy (agent0smith):

Change it to an improper fraction.

OpenStudy (agent0smith):

btw how did you get 132/99? That isn't correct at all.

OpenStudy (agent0smith):

Oh is it 8132/99? that...isn't remotely correct either...

OpenStudy (angelina__nievse):

My choices are A 8 132/99 B 8 132/999 C 8 1321/9999 D 8 999/132

OpenStudy (mjdennis):

Here is the way that I know how to solve: For example, pretend we are trying to find a/b for 0.901901901... It repeats every 3 digits, so multiply by 1 with 3 zeros, or 1000 1000*(a/b) = 901.901901... now subtract one (a/b) from each side... 1000*(a/b) - (a/b) = (901.901901...) - (0.901901901...) When you do the math, all the repeating part goes away: 999*(a/b) = 901 so (a/b) = 901/999 Good with the example?

OpenStudy (angelina__nievse):

So would it be B

OpenStudy (agent0smith):

Don't throw out a random letter answer. Show your work.

OpenStudy (mjdennis):

Don't know. Look at the discussion, you never gave any options.... and what @agent0smith said.

OpenStudy (agent0smith):

This question is made incredibly easy by it being multiple choice. Plug every option into a calculator until you find the right one?

OpenStudy (angelina__nievse):

Ok